# Topological Symmetry Breaking in Antagonistic Dynamics

### v2

@@
A dynamic concordia discors, a finely tuned equilibrium between opposing forces, is hypothesized to drive historical transformations 1. Similarly, a precise interplay of excitation and inhibition, the 80:20 ratio, is at the basis of the normal functionality of neural systems 2, 3, 4. In artificial neural networks, reinforcement learning allows for fine-tuning internal signed connections, optimizing adaptive responses to complex stimuli, and ensuring robust performance 5, 6, 7. At present, engineered structures of competing components are, however, largely unexplored, particularly because their emergent phases are closely linked with frustration mechanisms in the hosting network 8, 9. In this context, the spin glass theory has shown how an apparently uncontrollable non-ergodic chaotic behavior arises from the complex interplay of competing interactions and frustration among units 10, 11, leading to multiple metastable states preventing the system from exploring all accessible configurations over time 12. Here, we tackle the problem of disentangling topology and dynamics in systems with antagonistic interactions. We make use of the signed Laplacian operator to demonstrate how fundamental topological defects in lattices and networks percolate, shaping the geometrical arena and complex energy landscape of the system. This unveils novel, highly robust multistable phases and establishes deep connections with spin glasses when thermal noise is considered, providing a natural topological and algebraic description of their still-unknown set of pure states at zero temperature.
@@
Competing interactions introduce intrinsic incompatibilities between fundamental dynamical interactions and the elementary scales of the underlying lattice geometry 13, shaping the collective behavior of the system. This gives rise to new fundamental phenomena and yields to intriguing effects, such as the formation of exotic states like spin ice 14, 15, spin liquids 16, and spin glasses that have tantalized physicists for the last 40 years 17, 8. The latter, the paradigm of disordered magnetic systems, are characterized by competing interactions among their constituent spins, leading to an intricate and unpredictable dynamical and statistical configurational behavior 11. Signed interactions induce frustration and play a crucial role in describing the complex dynamics of these systems.
@@
The thorough study of spin glasses opened the door to one of the most elegant proposals in the field of modern theoretical physics: the replica trick and the replica symmetry breaking (RSB), which offers a comprehensive understanding of their non-ergodic thermodynamic properties 18, 19, 20 in the mean-field scenario. Solid efforts have been made to detect the nature of the spin glass phase 12, 10, 21. Therefore, significant progress in both analytical and numerical aspects has been made: the discovery of temperature and disorder chaos 22, stochastic stability 23, metastases 24, or the rigorous analysis of short-range spin glasses 25, 26. In a pioneering work, Amit et al. established the rigorous connection between neural networks and Ising spin glasses, highlighting that both systems share an energy landscape with numerous metastable states and valleys that enable the delocalized storage of patterns 24.
@@
However, several essential questions remain to be solved. Little is known about the number of pure ground states or the nature of broken symmetry in short-range spin glasses 27. Moreover, determining the ground states of Ising spin glasses in three or more dimensions is an NP-hard problem 28, which is closely related to many other hard combinatorial optimization problems 29. Similarly, disentangling new ensembles of correlated disorder 8, that is, predicting spin glass states in networks, represents an open challenge that remains crucial. In arbitrary architectures, in turn, it has never been fully understood the interplay between topology and signed interactions nor modeled such a delicate balance to generate multiple local states. Specifically, understanding the impact of competing units, i.e., elementary defects, in regular lattices and complex networks is key to properly controlling quantum error-correcting codes 30 and developing innovative brain-inspired devices, or even entirely new classes of structured artificial intelligence models that allow for more efficient and richer scenarios of machine-learning algorithms.
@@
Fig. LABEL:fig:a:lattc2sq_panel illustrates the SL in terms of spectral embedding arguments, as originally developed in 34 using the concept of antipodal proximity. Ordinary Laplacian embedding relies on the non-trivial solutions of the Laplace equation $L\bm{x}=0$ for placing vertices of a graph in Euclidean space. In this way, each vertex gets positioned at the weighted mean of its neighbors’ coordinates, with weights determined by the connecting edges. In contrast, when using $\bar{L}\bm{x}=0$ for embedding a signed graph, negatively connected neighbors contribute their coordinates with an opposite sign to the weighted sum determining the node placement. This makes nodes get closer to the antipodes of their oppositely tied neighbors. Moreover, since the SL has been formally proven to be a positive semidefinite operator (see Methods), we can rigorously implement thee recently introduced Laplacian framework 33, 32 to scan different spatial scales when an arbitrary architecture contains antagonistic or frustrated interactions. In particular, the heat capacity of a signed graph
@@
Fig. LABEL:fig:b:lattc2sq_panel shows a sketch of a 2D lattice with $M$ edges, where a random fraction of negative edges, denoted by $p=M^{-}/M$, is introduced, being $M^{-}$ the total number of negative edges. Fig. LABEL:fig:c:lattc2sq_panel shows the average heat capacity $\overline{C}$ for different values of $p$ in 2D squared lattices. In particular, we observe that the peak at short diffusion times disappears as the number of negative links increases, disrupting the lattice translational invariance. Note that the point where $\overline{C}$ exhibits a peak at short times formally corresponds to the so-called ultraviolet cutoff $\Lambda$, i.e., the lattice spacing 33, 36, 37. The existence of this peak, together with the subsequent plateau, is closely related to the existence of translational invariance in the network38, 37, which is lost at some specific value $p_{\rm c}$ when negative links gradually replace the positive ones. The probability distribution of the first maximum of $C$ for different realizations of 2D lattices with variable levels of frustration is reported in Fig. LABEL:fig:d:lattc2sq_panel, exhibiting bistable behavior and a Maxwell-like point for $p_{\rm c}\approx 0.1$. This symmetry breaking in the very topology plays, as discussed below, a key role in determining, altering, and disrupting dynamic phases and phase transitions.
@@
Figure 1: Signed Laplacian. (a) The antipodal proximity concept in the 2D spectral embedding with the SL 34. The position $\bm{x}_{i}$ of the $i^{\mathrm{th}}$ node is computed as the average position of the neighbors $\bm{x}_{j}$ weighted by the adjacency index $A_{ij}$. When $\operatorname{sign}(A_{ij})=-1$, the position $\bm{x}_{i}$ is updated in the direction opposite to $\bm{x}_{j}$. Black arrows correspond to coordinate axes. (b) Sketch of a 2D square lattice with random positive (blue links) and negative (red links) interactions. (c) Heat capacity, $C$, versus the temporal resolution parameter of the system, $\tau$, for different values of $p$ averaged over multiple disorder realizations (see legend). (d) Probability distribution, $-\log[P(\tau^{(0)}_{\max})]$, for the position, $\tau^{(0)}_{\max}$, of the first detected maximum of $C$ at varying $p$. Its shape changes from a single minimum for low $p$ to a bistable configuration at $p_{\rm c}$.
@@
Figure 2: Percolation of topological defects in regular structures. (a) Partitioning of a square 2D lattice as indicated by the positive and negative signs of $|\lambda_{0}\rangle.$ A set of fractal clusters emerges beyond the critical value, $p_{\rm c}$. (b) The lower part of each figure shows the partitioning of a 3D lattice as indicated by the positive and negative signs of $|\lambda_{0}\rangle$ (blue and red nodes, respectively). The upper part of the figure enhances only the positive ones. (c) Order parameter, $P_{\infty}$ versus the fraction of negative links, $p$ for a squared 2D lattice of variable size (see legend). Inset: Variance of the order parameter, $\chi$, as a function of the fraction of negative links, $p$. Note that $\chi$ diverges as the size of the system increases. Black dashed lines represent the expected value for the spin-glass transition at $T=0$. (d) Cluster size distribution for a 2D square lattice of different system sizes (see legend). The black dashed line represents the estimated Fisher exponent, $\tau\approx 1.8$. (e) Order parameter, $P_{\infty}$ versus the fraction of negative links, $p$, for a 3D cubic lattice of variable size (see legend). Inset: Variance of the order parameter, $\chi$, as a function of the fraction of negative links, $p$. Note that $\chi$ diverges as the size of the system increases. Black dashed lines represent the expected value for the spin-glass transition at $T=0$. (f) Cluster size distribution for a cubic lattice of different system sizes (see legend). The black dashed line represents the estimated Fisher exponent, $\tau=2.18$. (g) Sketch of topological defects in a squared lattice: single links (S), Z defects representing the elementary cell, and X defects representing “anti-nodes”. (h) $P_{\infty}$ versus the fraction of Z defects, $p$ for a squared 2D lattice of variable size (see legend). Inset: Variance of the order parameter, $\chi$, as a function of the fraction of Z defects, $p$. Black dashed vertical line corresponds to $p_{\rm c}=0.1$. (i) $P_{\infty}$ versus the fraction of X defects, $p$ for a squared 2D lattice of variable size (see legend). Inset: Variance of the order parameter, $\chi$, as a function of the fraction of X defects, $p$. Black dashed vertical line corresponds to $p_{\rm c}=0.1$.
@@
We still lack any indicator of a structural phase transition, as the observed fraction of negative links does not have an immediate interpretation in terms of site or bond percolation problems 39. A thorough analysis of the SL by exploring the parallel with the heat-like equation $\partial_{t}\psi=-\bar{L}\psi$ in the continuum limit approximation, leads to the expression $\bar{L}\equiv-\operatorname{\nabla^{2}}+B$, where $B$ is an arbitrary operator. This supports the analysis of any signed network as a perturbation to the usual network Laplacian, being formally analogous to the case of point-like particles diffusing in a uniform medium and being absorbed by a single sink (see Methods). In particular, it can be easily demonstrated that a negative link generates an “effective” influence area, affecting the surrounding region. It comes natural to speculate that the inclusion of single isolated negative links can be interpreted as a continuum percolation problem, where the influence area will depend on the specific underlying topology. Although this illustration does not resolve our original problem, it offers a valuable physical interpretation of the phenomena that occur as antagonistic interactions are progressively introduced to the system.
@@
We propose to use the components of the eigenvector ($|\lambda_{0}\rangle$) corresponding to the smallest eigenvalue $\lambda_{0}$ of $\bar{L}$ to define a universal order parameter for the emergence of a topological phase transition. Specifically, we take the signs of $|\lambda_{0}\rangle$ to identify independent mesoscopic clusters, with the most abundant elements assigned a positive sign. Note that in the limiting case $p=0$, $|\lambda_{0}\rangle$ is a uniform vector, but as $p$ increases, it develops a mix of positive and negative components, capturing how the system fragments into separate clusters. This transition naturally connects to percolation theory, offering a general framework to analyze it.
@@
Fig. LABEL:fig:c:percolation_panel shows the percolation phase transition for a 2D square lattice where a double-peak percolation transition emerges. There, the fraction of points belonging to the largest cluster, $P_{\infty}=n_{\infty}/N$, is the order parameter, with $N$ the total system size and $n_{\infty}$ the number of sites belonging to the percolating cluster. The intrinsic fluctuations, $\chi\equiv\sqrt{N}\sigma_{P_{\infty}}$, diverge at $p_{\rm c}$ as expected for usual second-order phase transitions 40. As reported in the inset of Fig. LABEL:fig:c:percolation_panel, the first peak in the susceptibility, located at $p_{\rm c}\sim 0.10(1)$, is fully compatible with the expected threshold for frustrated plaquettes and/or the emergence of the spin glass phase at zero temperature 41. Further examples considering triangular and hexagonal lattices are reported in SI, predicting the threshold where spin glass phases emerge 41, 42 (we will discuss this important point below). Moreover, Fig. LABEL:fig:e:percolation_panel shows the percolation phase transition for a 3D cubic lattice, where a double peak in the system susceptibility emerges, with $p_{\rm c}=0.22(1)$ 43 (4D analysis are reported in SI). The cluster-size distribution (see Fig. LABEL:fig:d:percolation_panel and Fig. LABEL:fig:f:percolation_panel) confirms that the phase transition effectively belongs to the ordinary percolation universality class, as $P(S)\sim S^{-\tau}$, with $\tau\approx 2.18(1)$ for the 3D case. The case of 2D lattices is more subtle, as we observe an exponent $\tau\approx 1.75(1)$, still compatible with previous results in small-to-mid systems in continuum percolation problems, where logarithmic corrections may play a key role in the convergence of the cluster size distribution for low dimensions 44.
@@
The double peak in the susceptibility represents a delicate point that must be discussed in terms of the geometrical properties of percolation in Euclidean spaces. Specifically, in three dimensions, two spanning clusters can geometrically coexist due to the natural emergence of lower-dimensional percolation structures. This is a natural consequence of the Alexander-Orbach conjecture 45: as critical percolation clusters present a universal spectral dimension $d_{S}=\nicefrac{{4}}{{3}}$ they can only coexist in dimension $d\geq 3$. In fact, we observe how the double-peak structure represents a finite-size effect in 2D lattices, which disappears in the infinite-size limit. Instead, in 3D, this phenomenon survives in the thermodynamic limit (a fact that is fully compatible with recent observations of the coexistence of spin glass and ferromagnetic order in spin glass 3D systems 46).
@@
We report in Fig. LABEL:fig:h:percolation_panel and Fig. LABEL:fig:i:percolation_panel the phase transition when, respectively, Z or X defects are included in a 2D lattice. In particular, for both cases, the percolation threshold is equivalent to those of S defects, as they exhibit a percolation phase transition for the same values of $p_{\rm c}$ (see also SI for further analyses on different lattices). The case of X defects generates an entirely new situation. In fact, in the squared 2D lattice, the double-peak structure does not vanish in the thermodynamic limit, as the second peak is directly related to the site-percolation problem. Note that X defects in the low-density regime are not expected to produce any closed loop, leading to no local frustration of interactions (the system is, then, effectively balanced, i.e., the fundamental eigenvalue is equal to zero). This leads to new fundamental dynamic implications that are explored below when X defects are considered together with a dynamic process.
@@
Figure 3: Ising dynamics. (a) Qualitative $T-p$ phase diagram for the 2D squared lattice with S defects. The lower part shows the energy levels corresponding to the first 100 eigenstates of $\bar{L}$ (see colorbar), for two values of $p$, respectively, in the ferromagnetic phase and the spin-glass one. The lowest eigenstate of $\bar{L}$ has been marked in red. Note that, inside the spin-glass phase, all eigenstates come closer in energy, and their ordering is lost at the ’onset of degeneracy’ at $p_{\rm c}$. (b) Qualitative $T-p$ phase diagram for X defects computed with the SL eigenstates through numerical simulations in a 2D square lattice of size $N=64^{2}$. The breakup of the ferroelectric phase translates into the emergence of a metastable one where the SL eigenvectors now define multiple degenerate stable states characterized by local order up to the paralectric/ferroelectric phase transition. At a certain $p$, stable minima becomes unstable, leading to a spin-glass phase. The lower part shows the energy levels corresponding to the first 100 eigenstates. In this case, the energy degeneracy is not observed before $p=0.4$, where the spin-glass phase emerges. (c) Annealing and sudden quenching initializing the system in a low eigenmode $|\lambda_{2}\rangle$ spin configuration on a 2D squared lattice with $p=0.25$ fraction of X defects. We heat the system up to $T\approx 1.5$, i.e., smaller but closer to the critical Ising temperature. After relaxation, the system is suddenly quenched, observing that, up to border effects, it does not escape from the local minima associated to $\operatorname{sign}(|\lambda_{2}\rangle)$. (d) Same as the previous case but using S defects with the lowest SL eigenmode at the $p_{\rm c}$. In this case, the system is locally stable but starts the attractor surfing at much lower temperatures. (e) Local magnetization order parameter, $m^{\rm GC}$, as a function of the temperature and the fraction of X defects for $N=64\times 64$. (f)- (g) Spin overlap, $\langle\lambda_{i}\,|\,\mathopen{}\sigma_{\infty}\rangle$, as a function of the temperature for different lattice sizes for (f) X defects, using $|\lambda_{2}\rangle$, and (g) S defects, using $|\lambda_{2}\rangle$. The critical temperature is indicated in the insets as a vertical dashed line. Note how $T_{\rm c}$ is significantly greater for the case of X defects, while, in contrast, the initial configuration loses its stability at much lower temperatures for S defects. All curves have been averaged over $n_{\rm avg}=1000$ independent realizations.
@@
$\mathscr{H}=-\sum_{(i,j)}J_{ij}\sigma_{i}\sigma_{j},$ (3)

### v3

@@
A dynamic concordia discors, a finely tuned equilibrium between opposing forces, is hypothesized to drive historical transformations 1. Similarly, a precise interplay of excitation and inhibition, the 80:20 ratio, is at the basis of the normal functionality of neural systems 2, 3, 4. In artificial neural networks, reinforcement learning allows for fine-tuning internal signed connections, optimizing adaptive responses to complex stimuli, and ensuring robust performance 5, 6, 7. Engineered systems with antagonistic interactions remain comparatively unexplored, particularly because their emergent phases are closely linked with frustration mechanisms in the hosting network 8, 9. In this context, the spin glass theory has shown how an apparently uncontrollable non-ergodic chaotic behavior arises from the complex interplay of competing interactions and frustration among units 10, 11, leading to multiple metastable states preventing the system from exploring all accessible configurations over time 12. Here, we show how topology constrains dynamics in systems with antagonistic interactions. We make use of the signed Laplacian operator to demonstrate how fundamental topological defects in lattices and networks percolate, shaping the geometrical arena and complex energy landscape of the system. This unveils novel, highly robust multistable phases and establishes deep connections with spin glasses when thermal noise is considered, providing a natural topological and algebraic description of emergent multistability and non-ergodicity in frustrated systems.
@@
Competing interactions introduce intrinsic incompatibilities between fundamental dynamical interactions and the elementary scales of the underlying lattice geometry 13, shaping the collective behavior of the system. This gives rise to new fundamental phenomena and leads to intriguing effects, such as the formation of exotic states like spin ice 14, 15, spin liquids 16, and spin glasses that have tantalized physicists for the last 40 years 17, 8. The latter, the paradigm of disordered magnetic systems, are characterized by competing interactions among their constituent spins, leading to an intricate and unpredictable dynamical and statistical configurational behavior 11. Signed interactions induce frustration and play a crucial role in describing the complex dynamics of these systems.
@@
The thorough study of spin glasses opened the door to one of the most elegant proposals in the field of modern theoretical physics: the replica trick and the replica symmetry breaking (RSB), which offers a comprehensive understanding of their non-ergodic thermodynamic properties 18, 19, 20 in the mean-field scenario. Solid efforts have been made to detect the nature of the spin glass phase 12, 10, 21. Therefore, significant progress in both analytical and numerical aspects has been made: the discovery of temperature and disorder chaos 22, stochastic stability 23, metastates 24, or the rigorous analysis of short-range spin glasses 25, 26. In a pioneering work, Amit et al. established the rigorous connection between neural networks and Ising spin glasses, highlighting that both systems share an energy landscape with numerous metastable states and valleys that enable the delocalized storage of patterns 24.
@@
However, several essential questions remain unanswered. Little is known about the number of pure ground states or the nature of broken symmetry in short-range spin glasses 27. Moreover, determining the ground states of Ising spin glasses in three or more dimensions is an NP-hard problem 28, which is closely related to many other hard combinatorial optimization problems 29. Similarly, disentangling new ensembles of correlated disorder 8, that is, predicting spin glass states in networks, represents an open challenge that remains crucial. In arbitrary architectures, in turn, it has never been fully understood the interplay between topology and signed interactions nor modeled such a delicate balance to generate multiple local states. Specifically, understanding the impact of competing units, i.e., elementary defects, in regular lattices and complex networks is key to properly controlling quantum error-correcting codes 30 and developing innovative brain-inspired devices, or even entirely new classes of structured artificial intelligence models that allow for more efficient and richer scenarios of machine-learning algorithms.
@@
Figure LABEL:fig:a:lattc2sq_panel illustrates the SL in terms of spectral embedding arguments, as originally developed in 34 using the concept of antipodal proximity. Ordinary Laplacian embedding relies on the non-trivial solutions of the Laplace equation $L\bm{x}=0$ for placing vertices of a graph in Euclidean space. In this way, each vertex gets positioned at the weighted mean of its neighbors’ coordinates, with weights determined by the connecting edges. In contrast, when using $\bar{L}\bm{x}=0$ for embedding a signed graph, negatively connected neighbors contribute their coordinates with an opposite sign to the weighted sum determining the node placement. This makes nodes get closer to the antipodes of their oppositely tied neighbors. Moreover, since the SL has been formally proven to be a positive semidefinite operator (see Methods), we can rigorously implement the recently introduced Laplacian framework 33, 32 to scan different spatial scales when an arbitrary architecture contains antagonistic or frustrated interactions. In particular, the heat capacity of a signed graph
@@
Figure LABEL:fig:b:lattc2sq_panel shows a sketch of a 2D lattice with $M$ edges, where a random fraction of negative edges, denoted by $p=M^{-}/M$, is introduced, being $M^{-}$ the total number of negative edges. Fig. LABEL:fig:c:lattc2sq_panel shows the average heat capacity $\overline{C}$ for different values of $p$ in 2D square lattices. In particular, we observe that the peak at short diffusion times disappears as the number of negative links increases, disrupting the lattice translational invariance. Note that the point where $\overline{C}$ exhibits a peak at short times corresponds to the shortest resolvable topological interaction scale (by analogy with an ultraviolet cutoff), which in regular lattices coincides with the lattice spacing 33, 36, 37. The existence of this peak, together with the subsequent plateau, is closely related to the existence of translational invariance in the network38, 37, which is lost at some specific value $p_{\rm c}$ when negative links gradually replace the positive ones. The probability distribution of the first maximum of $C$ for different realizations of 2D lattices with variable levels of frustration is reported in Fig. LABEL:fig:d:lattc2sq_panel, exhibiting bistable behavior and a Maxwell-like point for $p_{\rm c}\approx 0.1$. This symmetry breaking in the very topology plays, as discussed below, a key role in determining, altering, and disrupting dynamic phases and phase transitions.
@@
Figure 1: Signed Laplacian. (a) The antipodal proximity concept in the 2D spectral embedding with the SL 34. The position $\bm{x}_{i}$ of the $i^{\mathrm{th}}$ node is computed as the average position of the neighbors $\bm{x}_{j}$ weighted by the adjacency index $A_{ij}$. When $\operatorname{sign}(A_{ij})=-1$, the position $\bm{x}_{i}$ is updated in the direction opposite to $\bm{x}_{j}$. Black arrows correspond to coordinate axes. (b) Sketch of a 2D square lattice with random positive (blue links) and negative (red links) interactions. (c) Heat capacity, $C$, versus the temporal resolution parameter of the system, $\tau$, for different values of $p$ averaged over multiple disorder realizations (see legend). (d) Probability distribution, $\log[P(\tau^{(0)}_{\max})$], for the position, $\tau^{(0)}_{\max}$, of the first detected maximum of $C$ at varying pp. Its shape changes from a single minimum for low $p$ to a bistable configuration at $p_{\rm c}$.
@@
Figure 2: Percolation of topological defects in regular structures. (a) Partitioning of a square 2D lattice as indicated by the positive and negative signs of $|\lambda_{0}\rangle.$ A set of fractal clusters emerges beyond the critical value, $p_{\rm c}$. (b) The lower part of each figure shows the partitioning of a 3D lattice as indicated by the positive and negative signs of $|\lambda_{0}\rangle$ (blue and red nodes, respectively). The upper part of the figure enhances only the positive ones. (c) Order parameter, $P_{\infty}$ versus the fraction of negative links, $p$ for a squared 2D lattice of variable size (see legend). Inset: Variance of the order parameter, $\chi$, as a function of the fraction of negative links, pp. Note that $\chi$ diverges as the size of the system increases. Black dashed lines represent the expected value for the spin-glass transition at $T=0$. (d) Cluster size distribution for a 2D square lattice of different system sizes (see legend). The black dashed line represents the estimated Fisher exponent, $\tau\approx 1.8$. (e) Order parameter, $P_{\infty}$ versus the fraction of negative links, $p$, for a 3D cubic lattice of variable size (see legend). Inset: Variance of the order parameter, $\chi$, as a function of the fraction of negative links, pp. Note that $\chi$ diverges as the size of the system increases. Black dashed lines represent the expected value for the spin-glass transition at $T=0$. (f) Cluster size distribution for a cubic lattice of different system sizes (see legend). The black dashed line represents the estimated Fisher exponent, $\tau=2.18$. (g) Sketch of topological defects in a squared lattice: single links (S), Z defects representing the elementary cell, and X defects representing “anti-nodes”. (h) $P_{\infty}$ versus the fraction of Z defects, $p$ for a squared 2D lattice of variable size (see legend). Inset: Variance of the order parameter, $\chi$, as a function of the fraction of Z defects, pp. The black dashed vertical line corresponds to $p_{\rm c}=0.1$. (i) $P_{\infty}$ versus the fraction of X defects, $p$ for a squared 2D lattice of variable size (see legend). Inset: Variance of the order parameter, $\chi$, as a function of the fraction of X defects, pp. The black dashed vertical line corresponds to $p_{\rm c}=0.1$.
The heat capacity $C(\tau)$ serves as a spectral telescope: short times $\tau$ probe the high-energy (local) defects, while long times probe the low-energy (global) topological structure. Thus, anomalies in $C(\tau)$ allow us to detect the precise scale at which topological symmetry breaking occurs, providing a bridge to the structural analysis of defects that follows.
@@
We still lack any indicator of a structural phase transition, as the observed fraction of negative links does not have an immediate interpretation in terms of site or bond percolation problems 39. A complementary physical interpretation comes from the heat-like equation $\partial_{t}\psi=-\bar{L}\psi$: near the defect-percolation threshold, the diverging correlation length motivates an effective coarse-grained description in which $\bar{L}$ behaves as a Laplacian perturbed by localized sink-like terms, i.e. $\bar{L}\approx-\operatorname{\nabla^{2}}+B$, where $B$ encodes the coarse-grained effect of antagonistic links (see Methods and SI). In particular, it can be easily demonstrated that a negative link generates an “effective” influence area, affecting the surrounding region. It comes natural to speculate that the inclusion of single isolated negative links can be interpreted as a continuum percolation problem, where the influence area will depend on the specific underlying topology  [^1] We use the term “topology” in the graph-theoretic sense, referring to connectivity patterns and frustrated cycles in signed graphs, and not in the sense of topological invariants in condensed-matter physics.. Although this illustration does not resolve our original problem, it offers a valuable physical interpretation of the phenomena that occur as antagonistic interactions are progressively introduced to the system.
@@
We propose to use the components of the eigenvector ($|\lambda_{0}\rangle$) corresponding to the smallest eigenvalue $\lambda_{0}$ of $\bar{L}$ to define a universal order parameter for the emergence of a topological phase transition. Specifically, we take the signs of $|\lambda_{0}\rangle$ to identify independent mesoscopic clusters, with the most abundant elements assigned a positive sign. Note that in the limiting case $p=0$, $|\lambda_{0}\rangle$ is a uniform vector, but as $p$ increases, it develops a mix of positive and negative components, capturing how the system fragments into separate clusters. This transition naturally connects to percolation theory, offering a general framework to analyze it. The lowest eigenmode thus plays a dual role, capturing both structural fragmentation and the long relaxation bottlenecks induced by near-zero eigenvalues.
@@
Figure LABEL:fig:c:percolation_panel shows the percolation phase transition for a 2D square lattice where a double-peak percolation transition emerges. There, the fraction of points belonging to the largest cluster, $P_{\infty}=n_{\infty}/N$, is the order parameter, with $N$ the total system size and $n_{\infty}$ the number of sites belonging to the percolating cluster. The intrinsic fluctuations, $\chi\equiv\sqrt{N}\sigma_{P_{\infty}}$, diverge at $p_{\rm c}$ as expected for usual second-order phase transitions 41. As reported in the inset of Fig. LABEL:fig:c:percolation_panel, the first peak in the susceptibility, located at $p_{\rm c}\sim 0.10(1)$, is fully compatible with the expected threshold for frustrated plaquettes and/or the emergence of the spin glass phase at zero temperature 42. Further examples considering triangular and hexagonal lattices are reported in SI, predicting the threshold where spin glass phases emerge 42, 43 (we will discuss this important point below). Moreover, Fig. LABEL:fig:e:percolation_panel shows the percolation phase transition for a 3D cubic lattice, where a double peak in the system susceptibility emerges, with $p_{\rm c}=0.22(1)$ 44 (4D analyses are reported in SI). The cluster-size distribution (see Fig. LABEL:fig:d:percolation_panel and Fig. LABEL:fig:f:percolation_panel) confirms that the phase transition effectively belongs to the ordinary percolation universality class, as $P(S)\sim S^{-\tau}$, with $\tau\approx 2.18(1)$ for the 3D case. The case of 2D lattices is more subtle, as we observe an exponent $\tau\approx 1.75(1)$, still compatible with previous results in small-to-mid systems in continuum percolation problems, where logarithmic corrections may play a key role in the convergence of the cluster size distribution for low dimensions 45.
@@
The double peak in the susceptibility represents a delicate point that must be discussed in terms of the geometrical properties of percolation in Euclidean spaces. Specifically, in three dimensions, two spanning clusters can geometrically coexist due to the natural emergence of lower-dimensional percolation structures. This is a natural consequence of the Alexander-Orbach conjecture 46: as critical percolation clusters present a universal spectral dimension $d_{S}=\nicefrac{{4}}{{3}}$ they can only coexist in dimension $d\geq 3$. In fact, we observe how the double-peak structure represents a finite-size effect in 2D lattices, which disappears in the infinite-size limit. Instead, in 3D, this behavior is consistent with survival in the thermodynamic limit (see SI, and being compatible with observations of coexistence of spin glass and ferromagnetic order in 3D systems 47).
@@
We report in Fig. LABEL:fig:h:percolation_panel and Fig. LABEL:fig:i:percolation_panel the phase transition when, respectively, Z or X defects are included in a 2D lattice. In particular, for both cases, the percolation threshold is equivalent to those of S defects, as they exhibit a percolation phase transition for the same values of $p_{\rm c}$ (see also SI for further analyses on different lattices). The case of X defects generates an entirely new situation. In fact, in the squared 2D lattice, the double-peak structure does not vanish in the thermodynamic limit, as the second peak is directly related to the site-percolation problem. Note that X defects in the low-density regime are not expected to produce any closed loop, leading to no local frustration of interactions (the system is, then, effectively balanced, i.e., the fundamental eigenvalue is equal to zero). Balance, however, does not imply trivial dynamics, as mesoscale fragmentation and competing basins can still emerge. This leads to new fundamental dynamic implications that are explored below when X defects are considered together with a dynamic process.
@@
Figure 3: Ising dynamics. (a) Qualitative TpT-p phase diagram for the 2D squared lattice with S defects. The lower part shows the energy levels corresponding to the first 100 eigenstates of $\bar{L}$ (see colorbar), for two values of $p$, respectively, in the ferromagnetic phase and the spin-glass one. The lowest eigenstate of $\bar{L}$ has been marked in red. Note that, inside the spin-glass phase, all eigenstates come closer in energy, and their ordering is lost at the ’onset of degeneracy’ at $p_{\rm c}$. (b) Qualitative TpT-p phase diagram for X defects computed with the SL eigenstates through numerical simulations in a 2D square lattice of size $N=64^{2}$. The breakup of the ferroelectric phase translates into the emergence of a metastable one where the SL eigenvectors now define multiple degenerate stable states characterized by local order up to the paraelectric/ferroelectric phase transition. At a certain $p$, stable minima becomes unstable, leading to a spin-glass phase. The lower part shows the energy levels corresponding to the first 100 eigenstates. In this case, the energy degeneracy is not observed before $p=0.4$, where the spin-glass phase emerges. (c) Annealing and sudden quenching initializing the system in a low eigenmode $|\lambda_{2}\rangle$ spin configuration on a 2D squared lattice with $p=0.25$ fraction of X defects. We heat the system up to $T\approx 1.5$, i.e., smaller but closer to the critical Ising temperature. After relaxation, the system is suddenly quenched, observing that, up to border effects, it does not escape from the local minima associated to $\operatorname{sign}(|\lambda_{2}\rangle)$. (d) Same as the previous case but using S defects with the lowest SL eigenmode at the $p_{\rm c}$. In this case, the system is locally stable but starts the attractor surfing at much lower temperatures. (e) Local magnetization order parameter, $m^{\rm GC}$, as a function of the temperature and the fraction of X defects for $N=64\times 64$. (f)- (g) Spin overlap, $\langle\lambda_{i}\,|\,\mathopen{}\sigma_{\infty}\rangle$, as a function of the temperature for different lattice sizes for (f) X defects, using $|\lambda_{2}\rangle$, and (g) S defects, using $|\lambda_{2}\rangle$. The critical temperature is indicated in the insets as a vertical dashed line. Note how $T_{\rm c}$ is significantly greater for the case of X defects, while, in contrast, the initial configuration loses its stability at much lower temperatures for S defects. All curves have been averaged over $n_{\rm avg}=1000$ independent realizations.
@@
$\mathscr{H}=-\underset{\langle i,j\rangle}{\sum}J_{ij}\sigma_{i}\sigma_{j},$ (3)